# Gapfree graphs and powers of edge ideals with linear quotients

### v1

@@ 4. Pentagon together with vertices
Let $w_{n}=u_{s}e_{2}\in{\mathcal{A}}(e_{2})$ for some $s\leq r$. Note that $(b)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{2})=(b)$. If $w_{i}\in{\mathcal{A}}(e_{2})$, it follows from the induction hypothesis that there exists $j<n$ such that $w_{j}:w_{n}$ is a variable dividing $w_{i}:w_{n}$, as desired. Let $w_{i}\in{\mathcal{A}}(e_{1})$. If $b\nmid u_{s}$, then $w_{i}:w_{n}\in(b)$ since each element in ${\mathcal{A}}(e_{1})$ is divisible by $b$. Thus, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear and we are done. Suppose $b\mid u_{s}$. Then either $e_{4}$ or $e_{5}$ divides $u_{s}$ since $e_{1}\nshortmid u_{s}$. Since $e_{2}e_{4}=e_{1}e_{6}$, it is not possible to have $e_{4}\shortmid u_{s}$ So, $e_{5}\shortmid u_{s}$ where $u_{s}=w^{\prime}e_{5}$ for a generator $w^{\prime}\in I(G)^{N-2}$. In this case, it is important to note that $e_{3}\nshortmid u_{s}$ as $e_{3}e_{5}=e_{1}e_{7}$. Consider the elements $w_{j}=w^{\prime}e_{2}e_{1},w_{k}=w^{\prime}e_{6}e_{1}$ and $w_{l}=w^{\prime}e_{8}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a),(w_{k}:w_{n})=(x)$ and $(w_{l}:w_{n})=(z)$. This means $(a,b,x,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}.$
@@
Let $w_{n}=u_{s}e_{3}\in{\mathcal{A}}(e_{3})$ for some $s\leq r$. First recall that $e_{1},e_{2}\nshortmid u_{s}$ and also note that $e_{5}\nshortmid u_{s}$ since $e_{3}e_{5}=e_{1}e_{7}$. The case $w_{i}\in{\mathcal{A}}(e_{3})$ follows from the induction hypothesis as in the previous case. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup({\mathcal{A}}(e_{2})$. Then, $(b,p)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:w_{n})=(b)$ and $(u_{s}e_{2}:w_{n})=(p)$. If $b\nmid u_{s}$ and $p\nmid u_{s}$, then $(w_{i}:w_{n})\subseteq(b,p)$ implies that $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear. So, the claim holds for this case. If $b\mid u_{s}$, then $e_{4}\shortmid u_{s}$. Let $u_{s}=w^{\prime}e_{4}$ for a generator $w^{\prime}\in I(G)^{q-2}$. Consider $w_{j}=w^{\prime}e_{3}e_{1}$ and $w_{k}=e_{7}e_{1}$ in ${\mathcal{A}}(e_{1})$ such that $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(q)$. This means $(a,b,p,q)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$. One can verify that the same containment holds for the remaining case $p\mid u_{s}$ (which is split into two cases: $e_{6}\shortmid u_{s}$ or $e_{8}\shortmid u_{s}$).
@@
Let $w_{n}=u_{s}e_{4}\in{\mathcal{A}}(e_{4})$ for some $s\leq r$. First observe that $(a)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{4})=(a)$. The case $w_{i}\in{\mathcal{A}}(e_{4})$ follows from the induction hypothesis. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$. Note that this collection contains all generators of $I(G)^{q}$ that are divisible by $a$ and also $a\nmid w_{n}$. This means $w_{i}:w_{n}\in(a)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
@@
Let $w_{n}=u_{s}e_{5}\in{\mathcal{A}}(e_{5})$ for some $s\leq r$. Note that $(a,x)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{1}:u_{s}e_{5})=(a)$ and $(u_{s}e_{4}:u_{s}e_{5})=(x)$. The case $w_{i}\in{\mathcal{A}}(e_{5})$ follows from the induction hypothesis. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})\cup{\mathcal{A}}(e_{4})$. Note that each such $w_{i}$ is divisible by $a$ or $x$ and also $a\nmid w_{n}$. If $x\nmid w_{n}$, then $w_{i}:w_{n}\in(a,x)$ and the claim holds. If $x\mid w_{n}$, then either $e_{6}\shortmid u_{s}$ or $e_{7}\shortmid u_{s}$. We consider the $e_{6}\shortmid u_{s}$ case but the other one leads to the same conclusion. Let $u_{s}=w^{\prime}e_{6}$ for a generator $w^{\prime}\in I(G)^{q-2}$. Consider the elements $w_{j}=w^{\prime}e_{4}e_{2}$ and $w_{k}=w^{\prime}e_{6}e_{4}$ where $j,k<n$ where $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(x)$. This means $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear and the claim holds in this case.
[…]
Let $w_{n}=u_{s}e_{7}\in{\mathcal{A}}(e_{7})$ for some $s\leq r$. Notice that $(a,b,p)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{3}:u_{s}e_{7})=(a),(u_{s}e_{4}:u_{s}e_{7})=(b)$ and $(u_{s}e_{6}:u_{s}e_{7})=(p)$. The case $w_{i}\in{\mathcal{A}}(e_{7})$ follows from the induction hypothesis. For the remaining case, note that $w_{i}$ is divisible by either $a$ or $b$ or $p$ while $w_{n}$ is not divisible by neither $a$ or $b$. Furthermore, $w_{n}$ is not divisible by $p$ because if it does, then $e_{8}\shortmid u_{s}$ which implies that $e_{7}e_{8}=e_{6}e_{9}$ divides $w_{n}$, a contradiction. Thus, $w_{i}:w_{n}\in(a,b,p)$ and $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear. Hence, the claim holds for this case.
@@
Let $w_{n}=u_{s}e_{8}\in{\mathcal{A}}(e_{8})$ for some $s\leq r$. Notice that $(a,x)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{s}e_{2}:u_{s}e_{8})=(a)$ and $(u_{s}e_{6}:u_{s}e_{8})=(x)$. The case $w_{i}\in{\mathcal{A}}(e_{8})$ follows from the induction hypothesis. For the remaining case, notice that each $w_{i}$ is divisible by either $a,b$ or $x$ while $w_{n}$ is not divisible by any of these three variables. Thus, $w_{i}:w_{n}\in(a,x)$ which means $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
@@
For the final case, we have $w_{n}=e_{9}^{N}$. Observe that $(e_{9}^{N-1}e_{8}:e_{9}^{N})=(p)$ and $(e_{9}^{N-1}e_{5}:e_{9}^{N})=(p)$ and $(e_{9}^{N-1}e_{7}:e_{9}^{N})=(x)$. Since $\{b,p,x\}$ is a vertex cover of $G\setminus\{e_{9}\}$, we have $w_{i}:w_{n}\in(b,p,x)$ for each $w_{i}\neq w_{n}$.
Now, we introduce a peculiar concept of finite graphs.
@@
We now come to the first main theorem in the present paper.
@@
Given $n\geq 7$, there exists a CDCC graph $\Gamma_{n}$ with $n$ vertices for which all powers $I(\Gamma_{n})^{t}$ with $t\geq 2$ have linear quotients.
@@
Let $\Gamma_{7}$ denote the CDCC graph of Figure 4. Since a finite graph obtained by duplicating a vertex from a CDCC graph is again CDCC graph, the desired result follows from Propositions 2.2 and 4.3.
@@
While the matching number of the graph in Figure 4 is $3$; in general, the matching number of a CDCC graph can be arbitrarily large. Indeed, duplicating both endpoints of an edge increases the matching number without changing the existing induced subgraphs and gapfreeness.
5. Vertex duplication and linear quotients of powers
The techniques in the prior sections can be expanded to produce additional classes of graphs whose powers have linear quotients. In Section 2 it was shown that linear quotients are preserved under vertex duplication. There is a second way to define vertex duplication, where the vertex and its duplicate are connected via an edge. In this section, we show that for a gapfree graph, this type of vertex duplication preserves linear quotients as long as the duplication preserves gapfree. We first fix notation for the two types of duplication.
Notation 5.1.
Let $G$ be a gap free graph. Fix a vertex $x\in V(G)$. Define $G^{x}$ to be the graph obtained from $G$ by duplicating $x$ as in Definition2.1. Let
$V(G^{x})=V(G)\cup\{y\}$
for some $y\not\in V(G)$. Define $H^{x}$ to be the graph with vertex set $V(H^{x})=V(G^{x})$ and $E(H)=E(G^{x})\cup\{\{x,y\}\}$.
For any $W\subseteq V(H^{x})=V$ and any monomial $w\in I(H^{x})^{s}$, define $w_{W}$ to be the monomial formed by localizing $w$ at $W$. That is, if $w=\prod_{v_{i}\in V}v_{i}^{s_{i}}$, then $w_{W}=\prod_{v_{i}\in W}v_{i}^{s_{i}}$.
Next, we state some useful facts regarding the duplicated vertices. The proofs are straightforward from the definitions and so are not included.
Lemma 5.2.
Let $G$ be a gapfree graph. Using 5.1,
  • $N_{G^{x}}(x)=N_{G^{x}}(y)$.
  • $N_{H^{x}}[x]=N_{H^{x}}[y]$
  • $H^{x}$ is gap-free if and only if $V(H^{x})\setminus N_{H^{x}}[x]$ is an independent set.
We are now ready to prove the second main theorem in the present paper.
Theorem 5.3.
Let $G$ be a gapfree graph and fix $x\in V(G)$. Using 5.1, assume $H^{x}$ is gapfree and and $I(G)^{s}$ has linear quotients for some $s\geq 1$. Then $I(H^{x})^{s}$ has linear quotients.
Proof.
For any generator $w$ of $I(H^{x})^{s}$ define
$\mu(w)=\min\{i:\frac{w}{(xy)^{i}}\in Gens(I(G^{x})^{s-i})\}.$
Note that $\mu(w)=0$ if and only if $w$ is a minimal generator of $I(G^{x})^{s}$. Set $V=V(H^{x})$. Partition $V$ into sets:
$Z$ $=$ $\{x,y\}$
$A$ $=$ $N_{G}(x)$
$B$ $=$ $V(G)\setminus N_{G}[x]$
By 5.2, $B$ is an independent set. Fix an order $>$ on the elements of $B$. By Proposition 2.2 there is a linear quotients ordering $w_{1},w_{2},\dots,w_{q}$ on the minimal generators of $I(G^{x})^{s}$. We will extend this order to a linear quotients ordering $w_{1},\dots,w_{q},w_{q+1},\dots,w_{r}$ on the minimal generators of $I(H^{x})^{s}$. To simplify the notation, we adopt the convention that $w_{1}>w_{2}>\dots>w_{r}$. For generators $w$ and $w^{\prime}$ of $I(H^{x})^{s}$ not belonging to $I(G^{x})^{s}$, set $w>w^{\prime}$ if:
  1. (1)
    $\mu(w)<\mu(w^{\prime})$, or
  2. (2)
    $\mu(w)=\mu(w^{\prime})$ and $\deg(w_{Z})<\deg(w^{\prime}_{Z})$, or
  3. (3)
  4. (4)
    $\mu(w)=\mu(w^{\prime})$, $\deg(w_{Z})=\deg(w^{\prime}_{Z})$, $|\deg_{x}(w)-\deg_{y}(w)|=|\deg_{x}(w^{\prime})-\deg_{y}(w^{\prime})|$, and $w_{B}>w^{\prime}_{B}$ in the lexicographic ordering on $B$, using the order fixed above.
Let $w=w_{i}$ for some $i\geq 2$. We will show that $J=(w_{1},\dots,w_{i-1}):w_{i}$ is generated by variables. We may assume that $\mu(w)\neq 0$ by definition of the order and Proposition 2.2. Since $B$ is an independent set, the edges of $H^{x}$ can be partitioned as follows:
$E_{1}$ $=$ $\{xy\}$
$E_{2}$ $=$ $\{ab\mid a\in A,b\in B\}$
$E_{3}$ $=$ $\{az\mid a\in A,z\in Z\}$
$E_{4}$ $=$ $\{aa^{\prime}\mid a,a^{\prime}\in A\}.$
We make two key observations which will be used in the proofs.
  • (i)
    Let $w$ be a generator of $I(H^{x})^{s}$ with $\mu(w)\neq 0$. For any representation
    $w=(xy)^{\mu(w)}e_{1}\dots e_{s-\mu(w)}$
    for edges $e_{1},\dots,e_{s-\mu(w)}$, we have $e_{j}\notin E_{4}$ for all $j=1,\dots,s-\mu(w)$.
  • (ii)
    $A$ is a subset of the set of generators of $J$.
$w=(xy)^{\mu(w)-1}(xa_{1})(ya_{2})e_{2}\dots e_{s-\mu(w)}$
which is a contradiction to the minimality of $\mu(w)$.
Proof of (ii). Let $a\in A$. Then $u:=\frac{w}{xy}(xa)$ is a minimal generator of $I(H^{x})^{s}$ with $\mu(u)\leq\mu(w)-1$. Therefore $u>w$ and $u:w=a\in J$, proving the claim.
Let $v>w$. We will show that there is a variable generator of $J$ dividing $v:w$.
$\deg(v_{A})\geq s-\mu(v)>s-\mu(w)=\deg(w_{A}).$
$\deg(v_{A})=s-\mu(v)=s-\mu(w)=\deg(w_{A}).$
$2\gamma+m=\deg(v_{Z})<\deg(w_{Z})=2\gamma+m^{\prime}.$
Since $v_{A}=w_{A}$ and from the above inequality, there exists $a\mid(f_{1}\dots f_{k})_{A}$ such that $a\mid(g_{1}\dots g_{m^{\prime}})_{A}$. Without loss of generality, we may assume that $f_{1}=ab$ for some $b\in B$ and $g_{1}=az$ for some $z\in Z$. If $b\mid(v:w)$, set $w^{\prime}:=\frac{w}{g_{1}}f_{1}=\frac{w}{z}b$. Then $w^{\prime}>w$ with $w^{\prime}:w=b$, which completes the proof.
Case 2.2: Suppose that $\deg(v_{Z})=\deg(w_{Z})$.
Case 2.2-(i): Suppose that $|\deg_{x}(v)-\deg_{y}(v)|<|\deg_{x}(w)-\deg_{y}(w)|$. Since the right side of the inequality cannot be 00, without loss of generality, assume that $\deg_{x}(w)>\deg_{y}(w)$. Then since $\deg(v_{Z})=\deg(w_{Z})$, we have $\deg_{y}(v)>\deg_{y}(w)$ and $y\mid v:w$. Since $\deg_{x}(w)>\deg_{y}(w)$ there is an $a\in A$ with $e_{1}=ax$ in the representation of $w$. By definition of $H^{x}$, $a\in N(y)$ also. Set $w^{\prime}:=\frac{w}{ax}ay$. Then $w^{\prime}>w$ since $|\deg_{x}(w^{\prime})-\deg_{y}(w^{\prime})|<|\deg_{x}(w)-\deg_{y}(w)|$, and $w^{\prime}:w=y$.
Case 2.2-(ii): Suppose that $|\deg_{x}(v)-\deg_{y}(v)|=|\deg_{x}(w)-\deg_{y}(w)|$. Assume for a moment that $v_{Z}\neq w_{Z}$. Then without loss of generality, we may write
  • $\deg_{x}(v)=r_{1}=\deg_{y}(w)$
  • $\deg_{x}(w)=r_{2}=\deg_{y}(v)$
for some integers $r_{1}>r_{2}$. Then $w^{\prime}:=\frac{w}{y}x>w$ and $x=w^{\prime}:w$ divides $v:w$. Therefore, we may assume $v_{Z}=w_{Z}$. For the sake of clarity, we remind that the following statements are assumed at this point of the proof:
  • $1\leq\mu(v)=\mu(w)=\gamma$
  • $v_{A}=w_{A}$
  • $v_{Z}=w_{Z}$.
We fix a presentation $v=(xy)^{\gamma}f_{1}f_{2}\cdots f_{t}$ and $w=(xy)^{\gamma}g_{1}g_{2}\cdots g_{t}$. For simplicity, we assume that $f_{i_{1}}\dots f_{i_{m}}\neq g_{j_{1}}\dots g_{j_{m}}$ for all $m$-fold products with $1\leq m\leq t$ since otherwise the argument below is still valid recursively after removing the common factors.
Fix $p$ greatest in the ordering fixed on $B$ such that $p\mid(v:w).$ We claim that there is $u>w$ such that $u:w=p$. After rearranging the indices, we may write $f_{1}=a_{1}p$ for some $a_{1}\in A$.
Step 1: If there exists $\delta_{1}\in Z$ or $\delta_{1}\in B$ with $p>\delta_{1}$ such that $g_{1}=a_{1}\delta_{1}$ after some rearrangement of indices, set
$u=\frac{w}{g_{1}}f_{1}=\frac{w}{(a_{1}\delta_{1})}(a_{1}p)$
Then $u>w$ in the ordering since $\deg(u_{Z})<\deg(w_{Z})$ in the first case and $u_{B}>w_{B}$ in the latter one. Therefore, we may assume that $g_{1}=a_{1}\delta_{1}$ for some $\delta_{1}\in B$ with $\delta_{1}>p$.
Step 2: By the choice of $p$, we may set $f_{2}=a_{2}\delta_{1}$ for some $a_{2}\in A$. As before, if, up to reordering, $g_{2}=a_{2}\delta_{2}$ for some $\delta_{2}\in Z$ or $\delta_{2}\in B$ with $p>\delta_{2}$ we form the monomial
which satisfies the desired property. Therefore, we may assume that $\delta_{2}\in B$ and $\delta_{2}>p$.
Step 3: By the choice of $p$, we may set $f_{3}=a_{3}\delta_{2}$ for some $a_{3}\in A$. As before, if, up to reordering, $g_{3}=a_{3}\delta_{3}$ for some $\delta_{3}\in Z$ or $\delta_{3}\in B$ with $p>\delta_{3}$ we form the monomial
which satisfies the desired property. Therefore, we may assume $\delta_{3}\in B$ and $\delta_{3}>p$. The process must stop after $k$ steps when there is no $\delta_{k}\in B$ left with $\delta_{k}>p$.
Corollary 5.4.
Suppose that $I(G)^{s}$ has linear quotients for some $s$. Let $x$ be a vertex of $G$ such that non-neighbours of $x$ is an independent set. Let $H$ be the graph obtained from $G$ by replacing $x$ with a clique. Then $I(H)^{s}$ has linear quotients.
As an example of how this result can be applied, consider the class of graphs formed by replacing the central vertex $x$ of Figure 2 by a complete graph $K_{n}$, with each vertex of $K_{n}$ connected to $a,b,z$. We will denote such a graph by $(C_{5},K_{n})$ and will label the vertices of $K_{n}$ by $x_{1},\ldots,x_{n}$. The case $n=4$ is shown below.
$z$$p$$q$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$a$$b$
Figure 5. $(C_{5},K_{4})$
Example 5.5.
Let $I$ be the edge ideal of the graph $(C_{5},K_{n})$ for some $n\geq 1$. Then $I^{s}$ has linear quotients for all $s\geq 2$. The case $n=1$ holds by 4.2. For $n>1$ the result holds via induction using Proposition 2.2.
Notice that the above argument cannot extend to replace any vertex of the exterior pentagon with an edge. In particular, this would result in a finite graph that is not gapfree. However, the technique can be applied to the central vertex in Figure 3, for instance.
@@ 5. Vertex duplication and linear quotients of powers
In the present section, we are faced with a challenge on a slightly modified version of Conjecture 3.1.
@@
Let $G$ be a finite graph on the vertex set $V(G)$ and $E(G)$ the set of its edges. An admissible edge ordering of $G$ is a total order $\succ$ of the edges in $G$ with following property: for any $\{a,b\},\{c,d\}\in E(G)$ such that $\{a,b\}\cap\{c,d\}=\emptyset$ and $\{a,b\}\succ\{c,d\}$, we have either all edges incident to $a$ are larger than $\{c,d\}$ or all edges incident to $b$ are larger than $\{c,d\}$.
In what follows, unless there is a misunderstanding, we use the notation $ab$ instead of $\{a,b\}$ for an edge of a finite graph.
@@
By the induction hypothesis, there exists an admissible ordering of the edges in $G^{\prime}$. We shall order the edges in $G$ by $xx_{1}\succ\dots\succ xx_{s}\succ\text{ edges in }G^{\prime}$. We will show that this is an admissible ordering of the edges in $G$. Consider any two edges $ab>cd$ with $ab\cap cd=\emptyset$ in this ordering. Clearly, $cd$ cannot be incident to $x$.

### v2

@@ 5. Pentagon together with one vertex
CASE 1: Let $w_{n}=u_{\ell}e_{2}\in{\mathcal{A}}(e_{2})$ for some $\ell\leq r$. Note that $(b)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{1}:u_{\ell}e_{2})=(b)$. If $w_{i}\in{\mathcal{A}}(e_{2})$, it follows from the induction hypothesis that there exists $j<n$ such that $w_{j}:w_{n}$ is a variable dividing $w_{i}:w_{n}$, as desired. Let $w_{i}\in{\mathcal{A}}(e_{1})$. If $b\nmid u_{\ell}$, then $w_{i}:w_{n}\in(b)$ since each element in ${\mathcal{A}}(e_{1})$ is divisible by $b$. Thus, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear and we are done. Suppose $b\mid u_{\ell}$. Then either $e_{4}$ or $e_{5}$ divides $u_{\ell}$ since $e_{1}\nshortmid u_{\ell}$. Since $e_{2}e_{4}=e_{1}e_{6}$, it is not possible to have $e_{4}\shortmid u_{\ell}$ So, $e_{5}\shortmid u_{\ell}$ where $u_{\ell}=w^{\prime}e_{5}$ for a generator $w^{\prime}\in I(G)^{s-2}$. In this case, it is important to note that $e_{3}\nshortmid u_{\ell}$ as $e_{3}e_{5}=e_{1}e_{7}$. Consider the elements $w_{j}=w^{\prime}e_{2}e_{1},w_{k}=w^{\prime}e_{6}e_{1}$ and $w_{l}=w^{\prime}e_{8}e_{1}$ in ${\mathcal{A}}(e_{1})$ where $(w_{j}:w_{n})=(a),(w_{k}:w_{n})=(x)$ and $(w_{l}:w_{n})=(z)$. This means $(a,b,x,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}.$
@@
CASE 2: Let $w_{n}=u_{\ell}e_{3}\in{\mathcal{A}}(e_{3})$ for some $\ell\leq r$. First recall that $e_{1},e_{2}\nshortmid u_{\ell}$ and also note that $e_{5}\nshortmid u_{\ell}$ since $e_{3}e_{5}=e_{1}e_{7}$. The case $w_{i}\in{\mathcal{A}}(e_{3})$ follows from the induction hypothesis as in the previous case. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})$. Then, $(b,p)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{1}:w_{n})=(b)$ and $(u_{\ell}e_{2}:w_{n})=(p)$. If $b\nmid u_{\ell}$ and $p\nmid u_{\ell}$, then $(w_{i}:w_{n})\subseteq(b,p)$ implies that $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear. So, the claim holds for this case. If $b\mid u_{\ell}$, then $e_{4}\shortmid u_{\ell}$. Let $u_{\ell}=w^{\prime}e_{4}$ for a generator $w^{\prime}\in I(G)^{s-2}$. Consider $w_{j}=w^{\prime}e_{3}e_{1}$ and $w_{k}=e_{7}e_{1}$ in ${\mathcal{A}}(e_{1})$ such that $(w_{j}:w_{n})=(a)$ and $(w_{k}:w_{n})=(q)$. This means $(a,b,p,q)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$. One can verify that the same containment holds for the remaining case $p\mid u_{\ell}$ (which is split into two cases: $e_{6}\shortmid u_{\ell}$ or $e_{8}\shortmid u_{\ell}$).
@@
CASE 3: Let $w_{n}=u_{\ell}e_{4}\in{\mathcal{A}}(e_{4})$ for some $\ell\leq r$. First observe that $(a)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{1}:u_{\ell}e_{4})=(a)$. The case $w_{i}\in{\mathcal{A}}(e_{4})$ follows from the induction hypothesis. Let $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$. Note that this collection contains all generators of $I(G)^{q}$ that are divisible by $a$ and also $a\nmid w_{n}$. This means $w_{i}:w_{n}\in(a)$. Hence, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
@@
CASE 4: Let $w_{n}=u_{\ell}e_{5}\in{\mathcal{A}}(e_{5})$ for some $\ell\leq r$. Note that $(a,x)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{1}:u_{\ell}e_{5})=(a)$ and $(u_{\ell}e_{4}:u_{\ell}e_{5})=(x)$. The case $w_{i}\in{\mathcal{A}}(e_{5})$ follows from the induction hypothesis. If $w_{i}\in{\mathcal{A}}(e_{1})\cup{\mathcal{A}}(e_{2})\cup{\mathcal{A}}(e_{3})$, then $w_{i}:w_{n}\in(a)$ as in the previous case since each such $w_{i}$ is divisible by $a$ and also $a\nmid w_{n}$. Let $w_{i}\in{\mathcal{A}}(e_{4})$. If $x\nmid w_{n}$, then $w_{i}:w_{n}\in(x)$ and the claim holds. If $x\mid w_{n}$, then either $e_{6}\shortmid u_{\ell}$ or $e_{7}\shortmid u_{\ell}$. We consider the case $e_{6}\shortmid u_{\ell}$ and note that the other case follows similarly. Let $u_{\ell}=w^{\prime}e_{6}$ for a generator $w^{\prime}\in I(G)^{s-2}$. Consider the elements $w_{j}=w^{\prime}e_{4}e_{5}$ and $w_{k}=w^{\prime}e_{4}e_{9}$. Note that $j,k<n$ and $(w_{j}:w_{n})=(b)$ and $(w_{k}:w_{n})=(z)$. This means $(a,b,x,z)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$. As in the previous cases, let $w_{i}=e_{4}^{t_{4}}e_{5}^{t_{5}}\cdots e_{9}^{t_{9}}$ where $t_{4}>0$ and $w_{n}=e_{5}^{t^{\prime}_{5}}\cdots e_{9}^{t^{\prime}_{9}}$. If $w_{i}:w_{n}\notin(b,x,z)$, then considering degrees of these variables results with $t_{4}\leq 0$, a contradiction. Thus, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear.
@@
CASE 5: Let $w_{n}=u_{\ell}e_{6}\in{\mathcal{A}}(e_{6})$ for some $\ell\leq r$. Observe that $(a,b)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{2}:u_{\ell}e_{6})=(a)$ and $(u_{\ell}e_{4}:u_{\ell}e_{6})=(b)$. The case $w_{i}\in{\mathcal{A}}(e_{6})$ follows from the induction hypothesis. For the remaining case, note that each $w_{i}$ is divisible by $a$ or $b$ while $w_{n}$ is not divisible by neither $a$ nor $b$. This means $w_{i}:w_{n}\in(a,b)$. Thus, $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear in this case.
CASE 6: Let $w_{n}=u_{\ell}e_{7}\in{\mathcal{A}}(e_{7})$ for some $\ell\leq r$. Notice that $(a,b,p)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{3}:u_{\ell}e_{7})=(a),(u_{\ell}e_{4}:u_{\ell}e_{7})=(b)$ and $(u_{\ell}e_{6}:u_{\ell}e_{7})=(p)$. The case $w_{i}\in{\mathcal{A}}(e_{7})$ follows from the induction hypothesis. For the remaining case, note that $w_{i}$ is divisible by either $a$ or $b$ or $p$ while $w_{n}$ is not divisible by neither $a$ or $b$. Furthermore, $w_{n}$ is not divisible by $p$ because if $p\mid w_{n}$, then $e_{8}\shortmid u_{\ell}$ which implies that $e_{7}e_{8}=e_{6}e_{9}$ divides $w_{n}$, a contradiction. Thus, $w_{i}:w_{n}\in(a,b,p)$ and $(w_{1},\ldots,w_{n-1}):w_{n}$ is linear. Hence, the claim holds for this case.
CASE 7: Let $w_{n}=u_{\ell}e_{8}\in{\mathcal{A}}(e_{8})$ for some $\ell\leq r$. Notice that $(a,x)\subseteq(w_{1},\ldots,w_{n-1}):w_{n}$ since $(u_{\ell}e_{2}:u_{\ell}e_{8})=(a)$ and $(u_{\ell}e_{6}:u_{\ell}e_{8})=(x)$. The case $w_{i}\in{\mathcal{A}}(e_{8})$ follows from the induction hypothesis. For the remaining case, notice that each $w_{i}$ is divisible by either $a,b$ or $x$ while $w_{n}$ is not divisible by any of these three variables. Observe that $u_{\ell}=e_{8}^{t_{8}}e_{9}^{t_{9}}$. If $t_{9}\geq 1$, then $u_{\ell}=w^{\prime}e_{9}$ for a generator $w^{\prime}\in I(G)^{s-2}$ and $(w^{\prime}e_{8}e_{5}:w^{\prime}e_{8}e_{9})=(b)$. If $t_{9}=0$, then $w_{n}=e_{8}^{s}$. Notice that, for each $w_{i}$, we have $w_{i}:w_{n}\in(a,x,q)$ since $\{a,x,q\}$ is a vertex cover of $G\setminus\{e_{8}\}$. Consider the monomial $w_{j}=e_{8}^{s-1}e_{9}$ where $j<n$ by the definition of the order ${\mathcal{N}}^{(s)}$. Since $(w_{j}:w_{n})=(q)$, the claim holds.
@@
CASE 8: For the final case, we have $w_{n}=e_{9}^{s}$. Observe that $(e_{9}^{s-1}e_{8}:e_{9}^{s})=(p)$ and $(e_{9}^{s-1}e_{5}:e_{9}^{s})=(b)$ and $(e_{9}^{s-1}e_{7}:e_{9}^{s})=(x)$. Since $\{b,p,x\}$ is a vertex cover of $G\setminus\{e_{9}\}$, we have $w_{i}:w_{n}\in(b,p,x)$ for each $w_{i}\neq w_{n}$. ∎
Now, we introduce a particular new class of finite graphs.
@@
We now come to the next main theorem in the present paper.
@@
Given $n\geq 7$, there exists a CDCC graph $\Gamma_{n}$ with $n$ vertices for which all powers $I(\Gamma_{n})^{s}$ with $s\geq 2$ have linear quotients.
@@
Let $\Gamma_{7}$ denote the CDCC graph of Figure 6. Since a finite graph obtained by duplicating a vertex from a CDCC graph is again CDCC graph, the desired result follows from Propositions 2.3 and 5.4.                                                         ∎
@@
While the matching number of the graph in Figure 6 is $3$, in general, the matching number of a CDCC graph can be arbitrarily large. Indeed, duplicating both endpoints of an edge increases the matching number without changing the existing induced subgraphs and gapfreeness.
@@
In the present section, we consider a slightly modified version of Conjecture 4.1. We show that under an additional assumption on the ordering, if $I^{2},\,I^{5},\,I^{6}$ and $I^{7}$ have linear quotients, then $I^{q}$ has linear quotients for all $q\geq 2$. Before stating the result, we introduce the concept of an admissible order.
@@
Let $G$ be a finite graph on the vertex set $V(G)$ and $E(G)$ the set of its edges. An admissible edge ordering of $G$ is a total order $\succ$ of the edges in $G$ with following property: for any $\{a,b\},\{c,d\}\in E(G)$ such that $\{a,b\}\cap\{c,d\}=\emptyset$ and $\{a,b\}\succ\{c,d\}$, either all edges incident to $a$ are larger than $\{c,d\}$ or all edges incident to $b$ are larger than $\{c,d\}$.
Recall that, unless there is a misunderstanding, we use the notation $ab$ instead of $\{a,b\}$ for an edge of a finite graph. While not every order is admissible, such orders are ubiquitous in that every graph has one.
@@
By the induction hypothesis, there exists an admissible ordering of the edges in $G^{\prime}$. We shall order the edges in $G$ by $xx_{1}\succ\dots\succ xx_{s}\succ\text{ edges in }G^{\prime}$. We show that this is an admissible ordering of the edges in $G$. Consider any two edges $ab\succ cd$ with $ab\cap cd=\emptyset$ in this ordering. Clearly, $cd$ cannot be incident to $x$.